if f(x) be a polynomial of degree four such that f(-1)=-1, f(2)=-4,f(-3)=-9 and f(4)=-16 then find f(1)

Grade:12

1 Answers

Arun
5 years ago
Dear student
 Let
f(x) = x4 + ax³ + bx² + cx + d
Now
f(-1) = 1 -a +b -c +d = -1
f(-2) = 16 - 8a +4b -2c +d = -4
Similarly construct all the equations and find a,b,c,d and then form f(x)
And now you can find f(1)